OpenStudy (anonymous):

Find the distance between points to the nearest tenth. J(2,-1), K(2,5) Show all work.

OpenStudy (anonymous):

distance d= sqrt((2-2)^2 +(-1-5)^2) =sqrt(36) =6

OpenStudy (anonymous):

When I worked it out i got 3. isnt it y2 - y1..but did y1 - y2..so i think thats wrong.

OpenStudy (anonymous):

when you square it, it doesnt matter, whether it is y2-y1 or y1-y2, since (y2-y1)^2=(y1-y2)^2.

OpenStudy (anonymous):

y1=-1 y2=5 y1-y2=-6